Problem 7.62

A 63.0-kg skier starts from rest at the top of a ski slope of height 66.0m.

Part A

If frictional forces do ?1.08
Answer #1

Her energy at start is m*g*h, all gravitational potential energy. At the bottom, her net energy is m*g*h - Ef, (Ef is given as 1.08*10^4 J) and this is all kinetic energy = 0.5*m*v^2. Therefore

m*g*h - Ef = 0.5*m*v^2

v = ?[2*(m*g*h - Ef)/m]

On the horizontal path, the skier loses energy to two forces: friction Fr =